# Finding Safety Neurons in Large Language Models

### v1

@@ Abstract
Large language models (LLMs) excel in various capabilities but also pose safety risks such as generating harmful content and misinformation, even after safety alignment. In this paper, we explore the inner mechanisms of safety alignment from the perspective of mechanistic interpretability, focusing on identifying and analyzing safety neurons within LLMs that are responsible for safety behaviors. We propose generation-time activation contrasting to locate these neurons and dynamic activation patching to evaluate their causal effects. Experiments on multiple recent LLMs show that: (1) Safety neurons are sparse and effective. We can restore $90$% safety performance with intervention only on about $5$% of all the neurons. (2) Safety neurons encode transferable mechanisms. They exhibit consistent effectiveness on different red-teaming datasets. The finding of safety neurons also interprets “alignment tax”. We observe that the identified key neurons for safety and helpfulness significantly overlap, but they require different activation patterns of the shared neurons. Furthermore, we demonstrate an application of safety neurons in detecting unsafe outputs before generation. Our findings may promote further research on understanding LLM alignment. The source codes will be publicly released to facilitate future research.
Finding Safety Neurons in Large Language Models
Jianhui Chenthanks:  indicates equal contribution., Xiaozhi Wang11footnotemark: 1, Zijun Yao, Yushi Bai, Lei Hou, Juanzi Li Department of Computer Science and Technology, BNRist; KIRC, Institute for Artificial Intelligence, Tsinghua University, Beijing, 100084, China {chenjian22, wangxz20}@mails.tsinghua.edu.cn, {lijuanzi, houlei}@tsinghua.edu.cn
@@ 1 Introduction
In this work, we explore demystifying the mechanism behind safety alignment from the aspect of mechanistic interpretability (MI), which aims at reverse-engineering the neural models into human-understandable algorithms and concepts (Elhage et al., 2021). A typical MI pipeline includes attributing a model function to specific model components (e.g., neurons) and verifying that the localized components have causal effects on model behaviors with causal mediation analysis techniques like activation patching (Vig et al., 2020; Meng et al., 2022). However, existing MI methods mainly focus on interpreting tasks requiring only prompting (Wang et al., 2022a; Hanna et al., 2024) and few-token outputs (Dai et al., 2022; Wang et al., 2022b; Geiger et al., 2024). They cannot be directly applied to safety alignment, which requires model tuning and open-ended outputs. Previous work (Lee et al., 2024) interprets reducing toxicity as avoiding “toxicity vectors” in the generation, while this work tries to provide a holistic understanding of safety alignment beyond detoxification.
In this work, we propose a method for identifying safety-related neurons within LLMs (dubbed as safety neurons) and examining their causal effect on safety behaviors. Firstly, we introduce generation-time activation contrasting to calculate the change scores that quantify the importance of neurons to safety by comparing the neuron activations of the safety-aligned model and the unaligned model. We further propose dynamic activation patching to evaluate the causal effect of these neurons on the safety of long-range model generations, aiming to determine the minimal set of safety neurons that can effectively explain safety behaviors.
We investigate the effectiveness of the proposed method with three recent LLMs, including Llama-2 (Touvron et al., 2023), Mistral (Jiang et al., 2023), and Gemma (Team et al., 2024). Experiments show that we can consistently find safety neurons playing special roles in safety alignment with multiple desired properties: (1) Safety neurons are sparse and effective. We can restore $90$% safety performance with intervention only on about $5$% of all the neurons. (2) Safety neurons encode transferable mechanisms. Safety neurons are generally effective on multiple red-teaming benchmarks (Ji et al., 2024; Ganguli et al., 2022; Mazeika et al., 2024; Shen et al., 2023) without sacrificing general language modeling capability, which indicates they encode transferable safety mechanisms rather than shallow token filtering for specific datasets. (3) Safety neurons emerge stably. On different random trials, our method identifies essentially the same group of safety neurons.
@@
Moreover, safety neurons provide a potential explanation for the widely-recognized alignment tax issue (Askell et al., 2021; Ouyang et al., 2022). Specifically, the alignment tax here refers to the trade-off between harmlessness and helpfulness (Bai et al., 2022a), which means safety alignment enhances model safety (harmlessness) while sacrificing model capacity (helpfulness). We find that the key neurons identified by our method for safety and helpfulness have significant overlap, while the neurons found for other abilities like reasoning are distinct. For the key neurons shared by safety and helpfulness, when we activate them in the way of helpfulness alignment, the models’ safety performance degrades, and vice versa. This implies that alignment tax comes from requiring different activation patterns for the same neurons.
@@
To demonstrate the applications of safety neurons, we explore a straightforward case: LLM safeguard (Inan et al., 2023). We show that an effective unsafe generation detector can be built using the activations of safety neurons to predict, before actual generation, whether the response will contain harmful content. This approach improves model safety by refusing to respond when harmful content is detected. Specifically, the detector is a logistic regression model that uses the activations of top safety neurons as input. Experimental results show that adding this safeguard can significantly improve the safety of unaligned models and further enhance model safety after alignment.
To summarize, our contributions are three-fold: (1) We provide a mechanistic understanding of the safety alignment of LLMs by localizing key safety neurons and verifying their causal effect, transferability, and stability. (2) We interpret the alignment tax phenomenon by observing shared key neurons for helpfulness and harmlessness. (3) We explore the application of safety neurons in detecting unsafe generations. We hope the findings of safety neurons could facilitate future research on unveiling the inner workings of LLM alignment.
@@
Figure 1: The overview of our method. Neurons with significant activation differences between the aligned and unaligned models are identified using Generation-Time Activation Contrasting and assigned a change score. Dynamic Activation Patching then selects the number of neurons we need to have a strong enough causal effect on safety, referred to as safety neurons.
@@ 1.1 Safety Alignment
Due to the training instability and additional resources required by the reward model of RLHF, direct preference optimization (DPO) (Rafailov et al., 2024) has become a popular alternative (Tunstall et al., 2023; Ivison et al., 2023). The training efficiency can be further improved with minimal performance degeneration when combined with parameter-efficient fine-tuning methods (Sun et al., 2023; Hsu et al., 2024; Li et al., 2024). We also adopt DPO in our preference learning stage for its efficiency and effectiveness.
@@ 1.2 Neurons in Transformer
$h_{i}^{l+1}=h_{i}^{l}+\mathtt{MHA}^{l}(h_{i}^{l})+\mathtt{MLP}^{l}(h_{i}^{l}+\mathtt{MHA}^{l}(h_{i}^{l})).$
@@
$\mathtt{MLP}(x)=\mathrm{W_{down}^{\top}}(\sigma(\mathrm{W_{gate}}~{}x)\odot\mathrm{W_{up}}~{}x),$
@@
where $\mathrm{W_{down}},\mathrm{W_{gate}},\mathrm{W_{up}}\in\mathbb{R}^{d_{m}\times d}$ are projection matries, $\sigma(\cdot)$ is activation function, $\odot$ is element-wise product operator.
@@
MLP Neurons. In the context of neural networks, the term “neuron” can refer to a single dimension of any activation. We choose to study neurons in the intermediate layer of MLP (activation before down projection) since it has been shown such neurons encode meaningful and interpretable features (Wang et al., 2022b; Dai et al., 2022; Gurnee et al., 2023). Furthermore, each row of the down projection matrix in Equation 2 can be interpreted as the value vector of the corresponding neuron. This interpretation allows us to explore the semantics of neurons, as suggested by Geva et al. (2021).
2 Finding Safety Neurons
First, we introduce a general workflow of MI and discuss why it cannot be directly applied to interpret safety alignment. Then we introduce our method for locating safety neurons and evaluating their causal effects on safety alignment.
2.1 Mechanistic Interpretability Workflow
The first step in MI research typically involves identifying model components that have a critical impact on the targeted model function. Generally, this involves two steps. The first step is locating potential key model components (neurons, attention heads, etc.). For example, skill neurons (Wang et al., 2022b) are identified by calculating the predictivity on soft prompts; knowledge neurons (Dai et al., 2022) are identified through gradient attribution; directly enumerating all possible candidates (Wang et al., 2022a) is also adopted. The second step is to validate the causal effect of these identified components. Activation patching (Vig et al., 2020; Zhang and Nanda, 2023) is the most prevalent method for this purpose. In the model run with corrupted input prompts, the activation patching method patches the activations of investigated components with that on clean inputs and observes how much we can restore the probability or logits of predicting the next target token.
However, safety alignment involves open-ended generation, making previous methods, which are suitable only for tasks with a limited set of fixed target tokens, inapplicable. Enumerating all possible neuron group candidates is impractical for LLMs, while the sophisticated alignment problem cannot be expressed by input prompts. To address this, we propose generation-time activation contrasting to identify potential neuron candidates by contrasting the model activations before and after alignment. Furthermore, traditional activation patching typically intervenes only in the next token prediction, whereas safety evaluation requires long-form generation. We introduce dynamic activation patching to evaluate the causal effect of these neurons on the long-range dynamic generation process. The overview of our method is depicted in Figure 1. We first locate neurons with significant activation differences between the aligned and unaligned models using generation-time activation contrasting, followed by dynamic activation patching to determine the minimal set of neurons that have a strong enough causal effect on specific model behaviors.
2.2 Generation-Time Activation Contrasting
@@
We first introduce the method for identifying candidate neurons responsible for a specific ability in LLMs. Given two LLMs, $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, where $\mathcal{M}_{2}$ has acquired a specified ability through fine-tuning that $\mathcal{M}_{1}$ lacks, and this fine-tuning preserves the semantics of the components under investigation (for neurons, this refers to their corresponding key and value vectors introduced by Geva et al., 2021), such as through PEFT methods (Hu et al., 2021; Liu et al., 2022). For a given prompt $w=\langle w_{0},\ldots,w_{t}\rangle$, we denote the generation of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ as $w^{1}=\langle w_{t+1},\ldots,w_{t+m}\rangle$ and $w^{2}=\langle w^{\prime}_{t+1},\ldots,w^{\prime}_{t+n}\rangle$ respectively. The generation-time activation of $\mathcal{M}_{1}$ can be collected effectively with a forward pass on $\left[w,w^{1}\right]$ (the concatenation of prompt and generation, denoted as $\bar{w}^{1}$) and collect neuron activation on the token index from $t$ to $t+m-1$. The activation of $\mathcal{M}_{2}$ is also collected on $\bar{w}^{1}$ to ensure comparability of activations. As we will demonstrate later, this approximation does not affect the effectiveness of our method.
@@
Let $a^{(l)}_{i}\left(\mathcal{M}_{1};w\right)\left[j\right]\in\mathbb{R}$ be the activation of the $i$th neuron in layer $l$ of $\mathcal{M}_{1}$ at the $j$th token of a prompt $w$. Given the prompt dataset $\mathcal{D}$, we define the $\mathcal{M}_{1}$-based change score $\mathcal{S}^{(l)}_{i}(\mathcal{M}_{1},\mathcal{M}_{2};\mathcal{D})$ (and similarly for $\mathcal{M}_{2}$-based change score with the $\bar{w}^{1}$ replaced by $\bar{w}^{2}$ in the following equation) of $i$th neuron in layer $l$ as the root mean square of difference between generation-time activations of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$:
$\sqrt{\dfrac{\sum_{w\in\mathcal{D}}\sum_{j=|w|}^{\left|\bar{w}^{1}\right|-1}\left(a_{i}^{(l)}(\mathcal{M}_{1};\bar{w}^{1})[j]-a_{i}^{(l)}(\mathcal{M}_{2};\bar{w}^{1})[j]\right)^{2}}{\sum_{w\in\mathcal{D}}\left|w^{1}\right|}}.$
@@
To find safety neurons we choose the model after SFT as $\mathcal{M}_{1}$ (denoted as SFT) and the model after safety alignment as $\mathcal{M}_{2}$ (denoted as DPO). Then we sort all the neurons by the descending order of their change scores and use the top neurons as the safety neurons in experiments. Appendix D discusses some other potential design choices of our method.
@@ 2.3 Dynamic Activation Patching
To evaluate the causal effect of specific neurons in an open-ended generation scenario, we propose dynamic activation patching. This method involves a prompt, two models $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ (not necessarily the same as those in §2.2), and several forward passes: (1) Cache activations: run the model $\mathcal{M}_{2}$ on the current prompt and cache activations of given neurons; (2) Patched model run: run the model $\mathcal{M}_{1}$ on the same prompt with the activation of investigated neurons replaced by cached activation while the other neurons keep unchanged; (3) Get the next token prediction and append it to the prompt. Repeat these steps until finished.
To comprehensively evaluate the causal effect of safety neurons on LLMs’ safety behavior, in the experiments, $\mathcal{M}_{2}$ is DPO while $\mathcal{M}_{1}$ can be either SFT or the LLMs before SFT (denoted as Base).
3 Properties of Safety Neurons
In this section, we explore the properties (sparsity, effectiveness, transferability, and stability on training) of safety neurons with a series of experiments. The discussion of other properties of safety neurons can be found in appendix C.

### v2

# Towards Understanding Safety Alignment: A Mechanistic Perspective from Safety Neurons

@@ Abstract
Large language models (LLMs) excel in various capabilities but pose safety risks such as generating harmful content and misinformation, even after safety alignment. In this paper, we explore the inner mechanisms of safety alignment through the lens of mechanistic interpretability, focusing on identifying and analyzing safety neurons within LLMs that are responsible for safety behaviors. We propose inference-time activation contrasting to locate these neurons and dynamic activation patching to evaluate their causal effects on model safety. Experiments on multiple prevalent LLMs demonstrate that we can consistently identify about $5\%$ safety neurons, and by only patching their activations we can restore over $90\%$ of the safety performance across various red-teaming benchmarks without influencing general ability. The finding of safety neurons also helps explain the “alignment tax” phenomenon by revealing that the key neurons for model safety and helpfulness significantly overlap, yet they require different activation patterns for the same neurons. Furthermore, we demonstrate an application of our findings in safeguarding LLMs by detecting unsafe outputs before generation. The source code is available at https://github.com/THU-KEG/SafetyNeuron.
@@ 1 Introduction
In this work, we aim to demystify the mechanisms behind safety alignment from the perspective of mechanistic interpretability (MI), which focuses on reverse-engineering neural models into human-understandable algorithms and concepts (Elhage et al., 2021). A typical MI pipeline includes attributing model behaviors to specific model components and verifying that the localized components have causal effects on model behaviors with causal mediation analysis techniques like activation patching (Vig et al., 2020; Meng et al., 2022). However, existing MI methods (Wang et al., 2022a; Hanna et al., 2024; Geiger et al., 2024) mainly focus on attributing tasks requiring only prompting and few-token outputs to a limited search space of model components (e.g., attention heads). They cannot be directly applied to safety alignment, which naturally requires open-ended outputs and extensive model parameters as a high-level ability.
Considering that neurons are the most fundamental units in LLMs, and that prior work (Dai et al., 2022; Wang et al., 2022b; Gurnee et al., 2023, 2024) has shown that neurons encode a wide range of functionalities, we choose to investigate the safety mechanisms from a fine-grained, neuron-level perspective. We propose a two-stage framework (Figure 1) for identifying safety-related neurons (dubbed as safety neurons) and verifying their causal effects. The basic idea is that association is necessary for causality. Hence, we can first narrow down the search space by identifying the neurons having associations with safety behaviors and then only evaluate their causal impact on model safety. In the first stage, we employ inference-time activation contrasting to compute change scores, which quantify the association of neurons to safety. In the second stage, we propose dynamic activation patching to assess the causal effect of these neurons on the safety of long-range model outputs. Based on the framework, we make three-fold contributions:
  • We identify safety neurons across four open-sourced LLMs: Llama2-7B (Touvron et al., 2023), Mistral-7B (Jiang et al., 2023), Gemma-7B (Team et al., 2024), and Qwen2.5-3B (Qwen et al., 2025). We further demonstrate that: (1) Safety neurons identified by our framework are sparse and causally effective ($5\%$ of the neurons in unaligned models have over $90\%$ causal effects on safety alignment (§ 4.2), while the neurons identified via ablation-based method (Wei et al., 2024; Zhao et al., 2025) are not effective in our verification stage. (2) Safety neurons encode transferable mechanisms, which are generally effective on multiple red-teaming benchmarks without sacrificing generation quality (§ 4.3). (3) Safety neurons are robust to training randomness. In different random trials, our framework identifies essentially the same group of safety neurons (§ 4.4).
@@
  • We leverage safety neurons to provide a potential explanation for the widely-recognized alignment tax issue (Askell et al., 2021; Ouyang et al., 2022). Using our proposed framework, we find that the key neurons involved in the processes of safety alignment and helpfulness alignment exhibit significant overlap, while the neurons identified for other abilities like reasoning are less similar. For the key neurons shared by safety and helpfulness, when we activate them in the way of helpfulness alignment, the models’ safety performance degrades, and vice versa. This implies that alignment tax comes from requiring different activation patterns for a highly overlapping group of neurons (§ 5).
@@
  • We utilize safety neurons to develop an LLM safeguard (Inan et al., 2023), by showing that an effective unsafe generation detector can be built using the activations of safety neurons to predict, before actual generation, whether the response will contain harmful content. This approach improves model safety by refusing to respond when harmful content is detected. Experimental results show that adding this safeguard can significantly improve the safety of unaligned models and further enhance model safety after alignment (§ 6).
@@
Figure 1: Overview of the proposed framework. Neurons exhibiting significant activation differences between aligned and unaligned models are identified through inference-time activation contrasting and assigned a change score. Dynamic activation patching then selects the required number of neurons to achieve a strong causal effect on safety, referred to as safety neurons.
2 Preliminaries
@@ 2.1 Safety Alignment
Due to the training instability and additional resources required by the reward model of RLHF, direct preference optimization (DPO) (Rafailov et al., 2024) has become a popular alternative (Tunstall et al., 2023; Ivison et al., 2023). The training efficiency can be further improved with minimal performance degeneration when combined with parameter-efficient fine-tuning (PEFT) methods (Sun et al., 2023; Hsu et al., 2024; Li et al., 2024). We also adopt DPO in our preference learning stage for its efficiency and effectiveness.
@@ 2.2 Neurons in Transformer
$h_{i}^{l+1}=h_{i}^{l}+\mathtt{MHA}^{l}(h_{i}^{l})+\mathtt{MLP}^{l}(h_{i}^{l}+\mathtt{MHA}^{l}(h_{i}^{l})).$ (1)
@@
$\mathtt{MLP}(x)=\mathrm{W_{down}^{\top}}(\sigma(\mathrm{W_{gate}}~x)\odot\mathrm{W_{up}}~x),$ (2)
@@
where $\mathrm{W_{down}},\mathrm{W_{gate}},\mathrm{W_{up}}\in\mathbb{R}^{d_{m}\times d}$ are projection matrices, $\sigma(\cdot)$ is activation function, $\odot$ is element-wise product operator.
@@
MLP Neurons. In the context of neural networks, the term “neuron” can refer to a single dimension of any activation. We choose to study neurons in the intermediate layer of MLP (activation before down projection) since it has been shown such neurons encode diverse interpretable features (Wang et al., 2022b; Dai et al., 2022; Gurnee et al., 2023). Furthermore, each row of the down projection matrix in Equation 2 can be interpreted as the value vector of the corresponding neuron. This interpretation allows us to explore the tokens a neuron promotes or suppresses (Geva et al., 2021).
3 Finding Safety Neurons in LLMs
We first introduce how to locate neurons with significant activation differences between the aligned and unaligned models using inference-time activation contrasting, followed by dynamic activation patching to verify the causal effects on model behaviors and determine the minimal set of neurons.
3.1 Inference-time Activation Contrasting
@@
We first introduce the method for identifying candidate neurons responsible for the capabilities LLMs acquire through specific forms of training. Given two LLMs, $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, where $\mathcal{M}_{2}$ has acquired a specified ability through fine-tuning that $\mathcal{M}_{1}$ lacks, and this fine-tuning preserves the functionality of the components under investigation (for neurons, this refers to their corresponding key and value vectors introduced by Geva et al., 2021). For a given prompt $w=\langle w_{0},\ldots,w_{t}\rangle$, we denote the generation from $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ as $w^{1}=\langle w_{t+1},\ldots,w_{t+m}\rangle$ and $w^{2}=\langle w^{\prime}_{t+1},\ldots,w^{\prime}_{t+n}\rangle$ respectively. The inference-time activation of $\mathcal{M}_{1}$ can be collected effectively with a forward pass on $\left[w,w^{1}\right]$ (the concatenation of prompt and generation, denoted as $\bar{w}^{1}$) and collect neuron activation on the token index from $t$ to t+m1t+m-1. The activation of $\mathcal{M}_{2}$ is also collected on $\bar{w}^{1}$ to ensure comparability of activations. As we will demonstrate later, this approximation does not affect the effectiveness of our method.
@@
Let $a^{(l)}_{i}\left(\mathcal{M}_{1};w\right)\left[j\right]\in\mathbb{R}$ be the activation of the $i$th neuron in layer $l$ of $\mathcal{M}_{1}$ at the $j$th token of a prompt $w$, and denote the number of tokens in prompt $w$ as $\left|w\right|$. Given the prompt dataset $\mathcal{D}$, we define the $\mathcal{M}_{1}$-based change score $\mathcal{S}^{(l)}_{i}(\mathcal{M}_{1},\mathcal{M}_{2};\mathcal{D})$ (and similarly for $\mathcal{M}_{2}$-based change score with the $\bar{w}^{1}$ replaced by $\bar{w}^{2}$ in the following equation) of $i$th neuron in layer $l$ as the root mean square of difference between inference-time activations of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$:
$\sqrt{\frac{\sum_{w\in\mathcal{D}}\sum_{j=|w|}^{\left|\bar{w}^{1}\right|-1}\left(a_{i}^{(l)}(\mathcal{M}_{1};\bar{w}^{1})[j]-a_{i}^{(l)}(\mathcal{M}_{2};\bar{w}^{1})[j]\right)^{2}}{\sum_{w\in\mathcal{D}}\left|w^{1}\right|}}.$ (3)
@@
To find safety neurons we choose the model after SFT as $\mathcal{M}_{1}$ (denoted as SFT) and the model after safety alignment as $\mathcal{M}_{2}$ (denoted as DPO). Then we sort all the neurons by the descending order of their $\mathcal{M}_{1}$-based change scores computed on some safety-related datasets and use the top neurons as the safety neuron candidates in experiments. Appendix D discusses the difference between $\mathcal{M}_{1}$-based and $\mathcal{M}_{2}$-based change scores and some other potential design choices of our framework.
@@ 3.2 Dynamic Activation Patching
To evaluate the causal effect of specific neurons in an open-ended generation scenario, we propose dynamic activation patching. This method involves a prompt $w$, two models $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ (which may differ from the previous section), and several forward passes. Specifically, we repeat the following steps until the generation process is complete: (1) Cache activations: run the model $\mathcal{M}_{2}$ on the current prompt $w$ and cache the activations of the investigated neurons; (2) Patched model run: run the model $\mathcal{M}_{1}$ on the same prompt $w$ with the activation of investigated neurons replaced by cached activation while the other neurons keep unchanged; (3) Get the next token prediction and append it to the prompt $w$. A more detailed implementation can be found in Algorithm 1.
Let $\tilde{w}^{1}$ be the completed prompt obtained from dynamic activation patching, with all other notations consistent with those defined previously. Given the evaluation dataset $\mathcal{D}$, a metric $\mathcal{F}$ that assigns a real number score to each prompt, we define the causal effect $\mathcal{C}$ of specific neurons as follows:
$\mathcal{C}=\frac{\mathbb{E}_{w\in\mathcal{D}}\left[\mathcal{F}(\tilde{w}^{1})-\mathcal{F}(\bar{w}^{1})\right]}{\mathbb{E}_{w\in\mathcal{D}}\left[\mathcal{F}(\bar{w}^{2})-\mathcal{F}(\bar{w}^{1})\right]}.$ (4)
The intuition behind Equation 4 is that if specific neurons can faithfully explain the capabilities of model $\mathcal{M}_{2}$ that $\mathcal{M}_{1}$ lacks, then the causal effect $\mathcal{C}$ should be close to $1$. Conversely, a causal effect $\mathcal{C}$ close to 0 indicates a negligible causal effect.
To comprehensively evaluate the causal effect of safety neurons on LLMs’ safety behavior, we use DPO as $\mathcal{M}_{2}$, and $\mathcal{M}_{1}$ can be either SFT or the pre-trained LLMs before SFT (denoted as Base) in the following experiments unless otherwise specified.
4 Examining Safety Neurons
In this section, we explore the properties (sparsity, causal effect, transferability, and stability on training) of safety neurons with a series of experiments. The discussion of other properties of safety neurons can be found in Appendix C.
Figure 2: Causal effects of patching four models (both Base and SFT version) with activations from DPO, while applied on top safety neurons and random neurons, evaluated on Beavertails. The error bars are the 95% confidence interval over 55 random trials.